Find the mean, variance and standard deviation of the following measures: 2,5,6,3,9,10,12, and 15

1
Expert's answer
2022-03-17T03:40:07-0400

"\mu=\frac{1}{8}\Sigma_{i=1}^8x_i=\frac{62}{8}=7.75."

"\sigma^2=\frac{1}{8}\Sigma_{i=1}^8(x_i-\mu)^2=\frac{143.5}{8}=17.9375."

"\sigma=\sqrt{17.9375}\approx 4.2353."